OpenStudy (kkutie7):

Can someone explain this problem to me? Find the numbers b such that the average value of f(x) = 4 + 10x − 9x2 on the interval [0, b] is equal to 5.

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

\[\frac{f(b)-f(0)}{b-0}=5\] is the first step

OpenStudy (anonymous):

i didn't do anything yet really, just wrote down what the average value was now we plug some stuff in

OpenStudy (anonymous):

\[\frac{ 4 + 10b − 9b^2-4}{b}=5\]

OpenStudy (anonymous):

whao what happened there?

OpenStudy (dumbcow):

@satellite73 , isnt that for avg slope ? @Kkutie7 is this for calculus?

OpenStudy (kkutie7):

This is for calculus

OpenStudy (anonymous):

oh is this average value as in \[\frac{1}{b-0}\int_0^bf(x)dx=5\]?

OpenStudy (anonymous):

guess we can do that too thanks for the heads up @dumbcow

OpenStudy (kkutie7):

\[\frac{ 1 }{ b-a }\int\limits_{a}^{b}f(x)dx\]

OpenStudy (kkutie7):

ok I follow

OpenStudy (dumbcow):

:)

OpenStudy (anonymous):

\[\frac{1}{b}\int_0^b (4 + 10x − 9x^2)dx=5\]

OpenStudy (anonymous):

damn there are the question marks again is it just me?

OpenStudy (kkutie7):

nope I'm seeing it too

OpenStudy (kkutie7):

so I set the whole thing equal to 5?

OpenStudy (anonymous):

yeah and solve for \(b\)

OpenStudy (anonymous):

anti derivative is \(4x+5x^2-3x^3\) plug in \(b\) plug in \(0\) (which just gives zero) divide by \(b\) ,set it equal to 5 and solve for \(b\)

OpenStudy (anonymous):

mostly it is algebra

OpenStudy (kkutie7):

so first this right? \[\frac{ 1 }{ b }4x + 5x^{2} −3x^{3}, x=b\]

OpenStudy (anonymous):

i am getting those damned question marks again

OpenStudy (kkutie7):

so \[\frac{4b+5b^{2}-3b^{3}}{b}=5\]

OpenStudy (kkutie7):

\[4+5b-3b^2=5\] \[-3b^{2}+5b-1\] and quadratic?

OpenStudy (anonymous):

yeah but i would have put \[4+5b-3b^2=5\] solve the quadratic

OpenStudy (anonymous):

right, what you said

OpenStudy (anonymous):

\[3b^2-5b+1=0\] etc

OpenStudy (kkutie7):

b=.2324081, 1.4342585

OpenStudy (anonymous):

i guess

OpenStudy (kkutie7):

humm im not too sure

OpenStudy (anonymous):

why not?

OpenStudy (kkutie7):

Well I didn't get it right and I got a hint that reads: To find values b such that the average value of f on [0, b] is k, set up a function for fave with respect to b, and solve the equation fave(b) = k for b.

OpenStudy (anonymous):

isn't that what we did?

OpenStudy (kkutie7):

I thought so. the program I'm using is saying it's wrong

OpenStudy (anonymous):

does it want an exact answer, not a decimal?

OpenStudy (kkutie7):

probably... so Ill have to solve the quadratic by hand then

OpenStudy (anonymous):

you could use wolfram in you like

OpenStudy (anonymous):

Maybe mathway.com could help you? :)

OpenStudy (kkutie7):

well I can solve by hand it's just a pain.

OpenStudy (anonymous):

\[\frac{5\pm\sqrt{13}}{6}\] no charge

OpenStudy (kkutie7):

yeah i got the same i'm gonna see if the program excepts it

OpenStudy (kkutie7):

Yes it worked! Thank you =)

OpenStudy (anonymous):

yay! (never doubted it) yw